OpenStudy (stephaun):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). A:7 B:-14 C:0 D:14

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\] canyou put the points in that form where -4=x1 y1=8 x2=-3 y2=-6

OpenStudy (stephaun):

ok what next im a little confused

OpenStudy (stephaun):

Hello

OpenStudy (anonymous):

\[\frac{ -6-8 }{ -3+4 }\]

OpenStudy (anonymous):

can you simplify that

OpenStudy (stephaun):

-14 over 1

OpenStudy (stephaun):

So just -14

OpenStudy (anonymous):

now can you simplify \[-\frac{ 14 }{ 1 }\]

OpenStudy (anonymous):

yes

OpenStudy (stephaun):

ohh ok thanks

OpenStudy (anonymous):

do you understand how we did that?

OpenStudy (stephaun):

easier than expected

OpenStudy (stephaun):

yes

OpenStudy (anonymous):

yes when you're given 2 points and it asks for slope put it in \[\frac{ y2-y1 }{ x2-x1 }\]form this is slope-intercept

OpenStudy (stephaun):

? didn't you already do that

OpenStudy (anonymous):

yes i'm just telling you thats slope intercept form the points are always placed in (x1,y1)(x2,y2)

OpenStudy (stephaun):

ohh ok i learned that i was just over thinking the problem

OpenStudy (stephaun):

Thank you for your help :3

OpenStudy (anonymous):

you're welcome